# Shared Hardships Strengthen Bonds: Negative Shocks, Embeddedness and Employee Retention

### v1

@@ 5.2 Potential Mechanism: Sentiment Change after Issues
To explain the potential mechanisms for the turnover rate changes observed in the previous section, we employ sentiment analysis to examine drivers’ affective response. We use the VADER (Valence Aware Dictionary and sEntiment Reasoner) algorithm to compute sentiment scores, ranging -1 to 1 (the larger, more positive), for each communication note in the data. The VADER algorithm is a popular lexicon and rule-based sentiment analysis tool, which combines a sentiment lexicon–a list of words labeled as either positive or negative according to their semantic orientation–with certain rules (e.g., how to handle negation, modifier, etc) to compute sentiment scores. Compared to other sentiment analysis tools (e.g.,TextBlob), the VADER is specifically attuned to sentiments expressed in microblog-like contexts (e.g., social media) (Hutto, 2014), which is appropriate for our analysis given the form and style of the communication notes in our data. [^17] For sentiment scores, we particularly use the compound score in the VADER, which combines positive, neutral and negative scores calculated in the algorithm.
We regress the sentiment scores on the same set of regressors used in the duration models in Table 5, and plot the coefficient estimates and their confidence intervals, including the lag terms, in Figure 2.181818The full set of results is contained in Table C1, Appendix C. Note that the scale of y-axis in the second panel of Figure 2 is adjusted to show the changes in coefficients more clearly. Driver issues all have a negative effect on the sentiment of conversations right after the issues. Because it is a common practice to classify sentences as positive, neutral or negative if the VADER sentiment score is greater than 0.05, between 0.05 and -0,05, or less than -0.05, respectively,191919 https://github.com/cjhutto/vaderSentiment#about-the-scoring the sizes of the coefficients at period = 0 (i.e., the period issues were identified) implies a substantial negative effect of driver issues on conversation sentiment.
However, the pattern of how the negative sentiment changes over time is significantly different across issue types. In case of “Broken-equipment only,” the negative sentiment turns to positive after one period and the positiveness continues. On the other hand, the negative sentiment on “Misaligned only” remains negative. The coefficients on the combinations of the three issue types are less significant due to smaller numbers of such cases, but a similar pattern is observed in the second panel of Figure 2: as the case is more associated with “Aligned” issues, the negative effect tends to disappear faster. These observed relationships are consistent with companies taking remedial action in the event of broken equipment. It is plausible that the lagging positive sentiments are then the response to such actions.
@@ 6 Discussion
This finding is understandable within the framework of the unfolding model of labor turnover (Lee & Mitchell, 1994). Shocks are not direct determinants of retention, instead they highlight specific aspects of the current employment relationship in terms of links, fit, and sacrifice. Aligned shocks have the ultimate effect of underscoring job-fit compatibility between employee and employer. Motor carriers have the incentive to address Aligned issues immediately, leaving the employee feeling listened to and supported, thereby increasing embeddedness and ultimately retention. Indeed, sentiment analysis demonstrates positive employee feelings directly following equipment related shocks. Mis-aligned shocks have the opposite effect. Rather than highlighting job fit through interest alignment, these complaints underscore fundamental interest divergence at the current job.
@@
While pay is one of the most common complaints for truck drivers, its effect on retention is weak in comparison with home time, company and personal issues. While home time has been previously highlighted (Keller, 2002; Johnson et al., 2009, 2011), the company effects have not. This category includes policies with regards to pets, driver-facing cameras, and sleep apnea machines; however, the most common complaint in the category is the inability to take the tractor home at the end of a shift. Given that drivers’ homes, equipment yards and load locations can be separated by hundreds of miles, it is easy to understand the importance of this amenity. Managers may be willing to accept lower equipment utilization rates in favor of increased retention. Within the “personal” category the most important issue for drivers is being away from family. Offering more amenities in terms of family support is another important measure managers can focus on to increase retention.
The sheer variety of shock makes it impossible to entirely insulate employees. But negative shocks can have a silver lining when they underscore interest alignment. Our research uncovers a potentially under-used strategy for managers, which is to allow negative shocks to happen, but to address them immediately. In our context, the act of rescuing an employee from an equipment failure increases retention. Shared hardships strengthen the bonds of committment.
@@
Appendix A Fifteen Major Issue Categories and their Examples
  • Home-time: broken promises, shift change
  • Waiting: load/unload delays, waiting on assigned truck
  • Pay: inaccurate, not competitive
  • Equipment: broken/missing components, maintenance issues
  • Safety/Ethics: improperly loaded, unsafe operations
  • Training: too many hours, waiting on assigned truck
  • Team: co-driver issues, waiting on co-driver
  • Operations: fueling issues, unloading issues
  • Disrespect: slow/no response
  • Company: sleep apnea policy, pet policy
  • Covid: physically unable to work
  • Personal: family reasons
  • Health: health issues
  • Others: no show
Appendix B Robustness Check with Call Answer Rate
The same data and specifications in the models in Table 3 are used except for including an additional variable quantifying the responsiveness of each driver to the human resource company: the proportion of calls tagged with “voice mail,” “no answer,” etc. The log-pseudo-likelihood values of Model 1, 2, and 3 are -210,952.6, -211,002.9, and -184,545.8, respectively.
Table B1: Robustness Check with Call Answer Rate
Category Model 1 Model 2 Model 3 Category Model 1 Model 2 Model 3
Home-time 0.38 0.38 0.38 Operations -0.20 -0.20 -0.20
(s.e.) (0.03) (0.03) (0.03) (0.05) (0.05) (0.05)
[[[[t-stat]]]] [[[[10.89]]]] [[[[11.00]]]] [[[[10.92]]]] [[[[-4.25]]]] [[[[-4.22]]]] [[[[-4.26]]]]
{h.r.} {0.46} {0.46} {0.46} {-0.18} {-0.18} {-0.18}
Waiting -0.17 -0.16 -0.17 Disrespect 0.10 0.10 0.10
(0.04) (0.04) (0.04) (0.03) (0.03) (0.03)
[[[[-4.12]]]] [[[[-3.86]]]] [[[[-3.97]]]] [[[[2.98]]]] [[[[3.00]]]] [[[[2.97]]]]
{-0.16} {-0.15} {-0.15} {0.11} {0.11} {0.11}
Pay 0.22 0.22 0.22 Company 0.27 0.26 0.28
(0.02) (0.02) (0.02) (0.09) (0.09) (0.09)
[[[[9.65]]]] [[[[9.44]]]] [[[[9.71]]]] [[[[3.04]]]] [[[[2.94]]]] [[[[3.25]]]]
{0.25} {0.24} {0.25} {0.31} {0.30} {0.33}
Equipment -0.25 -0.24 -0.25 Covid 0.44 0.44 0.42
(0.02) (0.02) (0.02) (0.13) (0.13) (0.14)
[[[[-11.73]]]] [[[[-11.48]]]] [[[[-11.70]]]] [[[[3.26]]]] [[[[3.26]]]] [[[[3.12]]]]
{-0.22} {-0.22} {-0.22} {0.55} {0.55} {0.53}
Safety/Ethics 0.22 0.23 0.21 Personal 0.60 0.59 0.60
(0.05) (0.05) (0.05) (0.04) (0.04) (0.04)
[[[[4.67]]]] [[[[4.96]]]] [[[[4.52]]]] [[[[14.94]]]] [[[[14.74]]]] [[[[14.97]]]]
{0.25} {0.26} {0.24} {0.82} {0.81} {0.83}
Training 0.24 0.23 0.24 Health 0.32 0.32 0.31
(0.04) (0.04) (0.04) (0.06) (0.06) (0.06)
[[[[5.43]]]] [[[[5.13]]]] [[[[5.32]]]] [[[[5.26]]]] [[[[5.40]]]] [[[[5.01]]]]
{0.27} {0.26} {0.27} {0.37} {0.38} {0.36}
Team 0.08 0.07 0.09 Others 1.08 1.06 1.16
(0.09) (0.09) (0.09) (0.21) (0.19) (0.18)
[[[[0.93]]]] [[[[0.80]]]] [[[[1.00]]]] [[[[5.23]]]] [[[[5.47]]]] [[[[6.35]]]]
{0.08} {0.07} {0.09} {1.94} {1.89} {2.17}
Appendix C The Full Sentiment Regression Result
The outcome variable is the VADER (Valence Aware Dictionary and sEntiment Reasoner) score. The regression models include the same sets of covariates and fixed effects in the models in Table 5. Standard errors are clustered at the spell level.
Table C1: Change of Conversation Sentiment after Driver Issues

### v2

@@ 5.3 Potential Mechanism: Sentiment Change after Issues
To explain the potential mechanisms for the turnover rate changes observed in the previous section, we employ sentiment analysis to examine drivers’ affective response. We use the VADER (Valence Aware Dictionary and sEntiment Reasoner) algorithm to compute sentiment scores, ranging -1 to 1 (the larger, more positive), for each communication note in the data. The VADER algorithm is a popular lexicon and rule-based sentiment analysis tool, which combines a sentiment lexicon–a list of words labeled as either positive or negative according to their semantic orientation–with certain rules (e.g., how to handle negation, modifier, etc) to compute sentiment scores. Compared to other sentiment analysis tools (e.g.,TextBlob), the VADER is specifically attuned to sentiments expressed in microblog-like contexts (e.g., social media) (Hutto & Gilbert, 2014), which is appropriate for our analysis given the form and style of the communication notes in our data. [^22] For sentiment scores, we use the compound score in the VADER, which combines positive, neutral and negative scores calculated in the algorithm.
We regress the sentiment scores on the same set of regressors used in the Cox PH models in Table 6, and plot the coefficient estimates and their confidence intervals, including the lag terms, in Figure 3.232323The full set of results is contained in Table A1, Appendix A. Note that the scale of y-axis in the second panel of Figure 3 is adjusted to show the changes in coefficients more clearly. Driver issues all have a negative effect on the sentiment of conversations right after the issues. Because it is a common practice to classify sentences as positive, neutral or negative if the VADER sentiment score is greater than 0.05, between 0.05 and -0,05, or less than -0.05, respectively,242424 https://github.com/cjhutto/vaderSentiment#about-the-scoring the sizes of the coefficients at period = 0 (i.e., the period issues were identified) implies a substantial negative effect of driver issues on conversation sentiment.
However, the pattern of how the negative sentiment changes over time is significantly different across issue types. In case of “Broken-equipment only,” the negative sentiment turns to positive after one period and the positiveness continues. On the other hand, the negative sentiment on “Misaligned only” remains negative. The coefficients on the combinations of the three issue types are less significant due to the smaller number of such cases, but a similar pattern is observed in the second panel of Figure 3: when the case is more associated with “Aligned” issues, the negative effect tends to disappear faster. These observed relationships are suggestive evidence that companies take remedial action in the event of broken equipment. The lagging positive sentiments are then the response to such actions.
@@ 6 Discussion
This finding is consistent with our hypothesis and aligns well with the framework of the unfolding model of labor turnover (Lee & Mitchell, 1994). Shocks are not direct determinants of retention, instead they highlight specific aspects of the current employment relationship in terms of links, fit, and sacrifice. Aligned shocks have the ultimate effect of underscoring job-fit compatibility between employee and employer. Motor carriers have the incentive to address Aligned issues immediately, leaving the employee feeling listened to and supported, thereby increasing embeddedness and ultimately retention. Indeed, sentiment analysis demonstrates positive employee feelings directly following equipment-related shocks. Misaligned shocks have the opposite effect. Rather than highlighting job fit through interest alignment, these complaints underscore fundamental interest divergence at the current job.
The primary threat to this interpretation is self-reporting bias (Scott & Balthrop, 2021), for instance, if unobserved employee attachment is correlated with observed communications. If employees that are particularly attached to an organization are more likely report problems, then communication is the outcome of employee attachment, rather than the cause. Our extensive data set–the largest and richest panel yet used in trucking retention research– allows us to address these concern using three strategies: i) we control for non-response and complaint frequencies, ii) we restrict analysis to the sub-sample of observed responders, and, iii) we examine whether our results change by equipment ownership status in a theoretically consistent way. The main result stands in each case. Negative shocks increase retention when they highlight job-fit through employee-firm interest alignment.
6.1 Theoretical Implications
Our results have implications to retention theory generally, but also provide context-specific implications to retention research within the trucking sector. First, our results highlight the usefulness of the unfolding model of turnover. Our paper finds context-specific shocks to be resoundingly predictive of employment duration, consistent with the unfolding model. It would be difficult to model and understand these correlations within the framework of competing theories, viz., expected utility theory.
Second, our research highlights important boundary conditions within the unfolding model of labor turnover on how shocks translate into retention. When negative shocks underscore job-fit, the duration of employment increases. In the trucking industry context, it is equipment, operations, and wait-time shocks. At the boundary of retention lies the complementary theory of job embeddedness, which rests on job-fit, links and sacrifice. Given our findings on how job-fit influences the ultimate effect of shocks, it is reasonable to suspect that considerations of links and sacrifice may have similar effects, at least in other industries where these concerns are more significant.
These results also have intriguing links to an analogous stream of defensive marketing research showing how businesses are better served by devoting resources towards addressing existing customer complaints, than on offensive marketing campaigns designed to draw in new customers (Fornell & Wernerfelt, 1987). For example, by addressing service failures expeditiously, business can increase customer loyalty (Tax et al., 1998). Similarly, in high turnover-industries, employers may be better-served through defensive expenditures on current employees rather than devoting more effort to recruitment. Our data is produced by exactly this sort of defensive investment. A key point for theory to address given sophisticated retention efforts, then, is the role of employers in “voluntary” exits, which are generally ignored in the psychology literature: even these exits are to some extent joint decisions between employee and firm, potentially revealing a firm’s exhaustion with an employee.
6.2 Guidance for Practitioners
@@
While pay is one of the most common complaints for truck drivers, its effect on retention is weak in comparison with home time, company and personal issues. While home time has been previously highlighted (Keller, 2002; Johnson et al., 2009, 2011), the company effects have not. This category includes policies with regards to pets, driver-facing cameras, and sleep apnea machines; however, the most common complaint in the category is firm policies that do not allow a driver to take the tractor home at the end of a shift. Given that drivers’ homes, equipment yards and load locations can be separated by hundreds of miles, it is easy to understand the importance of this amenity. Managers may be willing to accept lower equipment utilization rates in favor of increased retention. Within the “personal” category the most important issue for drivers is being away from family. Offering more amenities in terms of family support is another important measure managers can focus on to increase retention.
To provide more general guidance, in Figure 4, we plot issues by relative frequency and employment impact (based on the coefficients in Table 3), where hollow squares plot the issues that increase the duration of employment and solid black points plot the issues that decrease the duration. We can classify issues into High Frequency & Medium Impact (Pay, Equipment), Low Frequency & High Impact (Personal, Covid, Home time, Health, Company) and Low Frequency & Low Impact (Training, Safety/Ethics, Operations, Waiting, Team, and Disrespect). Issues in the high-impact category are mostly idiosyncratic to drivers, company issues being a notable exception. On the other hand, equipment issues stand out as the highest frequency event, and an event that increases employment duration. Timely response to equipment-related concerns may be the best way for companies to reduce turnover.
[$=$ Figure 4 here $=$]
6.3 Future Research Directions
As with all archival research, there is potential for unobserved variables to confound estimates. In this study, the key worry is that unobserved driver type is correlated with communication frequency. While we have undertaken extensive robustness checks to ensure the reliability of our results, the ideal means to fully address the potential for confounding variation would be through a series of field experiments.
Based on our results, companies in high-turnover industries may consider performing experiments with on-boarding and orientation programs. These programs need not be entirely pleasant experiences, so long as they highlight job-embeddedness characteristics, and perhaps build group unity and identity. Randomly allocating employees to such experiments would allow researchers to isolate any confounding effects while examining the main effects that we find in this paper - shared hardships strengthen the bonds of commitment.
@@ 6.3 Future Research Directions
Appendix A The Full Sentiment Regression Result
The outcome variable is the VADER (Valence Aware Dictionary and sEntiment Reasoner) scores for the communication records. The regressions include the same sets of covariates and fixed effects as in the models in Table 6. Standard errors are clustered at the spell level.
Table A1: Change of Conversation Sentiment after Driver Issues